OpenStudy (anonymous):

imagine you are the owner of a pizza shop. That offers three different types of crust, seven different toppings and three different cheese. I NEED TO KNOW HOW MANY DIFFERENT COMBINATIONS THERE ARE/HOW TO GET THE ANSWER. i don't want the answer i just want to know how to find the answer of how many combinations there are.

jimthompson5910 (jim_thompson5910):

so you're only allowed one type of crust, one type of cheese, and one topping?

OpenStudy (alekos):

Just draw a simple tree diagram and you will see how many combinations there'll be

jimthompson5910 (jim_thompson5910):

If one of each type is allowed (1 crust, 1 cheese, 1 topping), then \[\large \text{Answer} = \begin{pmatrix} \text{Number}\\ \text{of}\\ \text{Crusts} \end{pmatrix} \times \begin{pmatrix} \text{Number}\\ \text{of}\\ \text{Cheeses} \end{pmatrix} \times \begin{pmatrix} \text{Number}\\ \text{of}\\ \text{Toppings} \end{pmatrix} \] I'm using the counting principle http://www.mathsisfun.com/data/basic-counting-principle.html